# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ 6.1.2. Removal
For natural language data, stopwords are often removed due to their high frequency and little meaningful contribution (e.g., ”a” and ”the”) (Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017). For code-formatted data, heuristic rules can be designed to clean the data effectively. Irrelevant elements, such as logical operators (Matmti and Fard, 2023), trailing white spaces (Matmti and Fard, 2023), punctuation marks (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021a; Bansal et al., 2023a; Koyuncu et al., 2019; Matmti and Fard, 2023; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017), comments (Zügner et al., 2021), empty lines (Zügner et al., 2021), and non-word characters (Bansal et al., 2023a), can also be removed to simplify the data and reduce noise.
@@ 6.1.3. Unifying
To organize natural language formatted data, common text unifying methods include using lowercase (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021a) and applying stemming (Koyuncu et al., 2019; Bangcharoensap et al., 2012; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017). These works simplify the data by standardizing variations in word form. Lowercasing ensures that all tokens are converted to lowercase while stemming groups of words with the same core meaning by retaining the root form of words. Both methods help models focus on the core meaning of the data rather than extra information.
@@
In addition to organizing data, handling sensitive information is crucial to ensure data purity and model security. Researchers replace numbers and strings in source code with unique tokens such as ¡num¿ (Nashid et al., 2023; Xu et al., 2022; Wang et al., 2021a; Zügner et al., 2021; Lin et al., 2021) and ¡str¿ (Nashid et al., 2023; Xu et al., 2022; Zhou et al., 2019; Wang et al., 2021a; Bansal et al., 2023a; Zügner et al., 2021; Lin et al., 2021). Numbers and strings can vary in length and format, requiring additional handling methods. Masking specific values encourages models to focus on the type of data. Besides, replacing them with unique tokens also helps protect private data, as numbers and strings may sometimes contain sensitive information.
@@ 6.1.4. Other
One of the preprocessing methods involves using ASTs to extract information from source code. In this case, ASTs are not considered part of the context but are used to extract semantic code knowledge from repositories (Bi et al., 2024b; Liao et al., 2024). This narrows the scope of source code context while facilitating retrieval by constructing semantic graphs (Phan et al., 2024).
Besides, researchers also use static analysis to convert source code into XML format (Zhou et al., 2024) or apply program slicing (Zhang et al., 2024a), enabling further analysis and processing. In addition, manual expertise (Zhao et al., 2024) and rule-based methods (Wang et al., 2023) are also applied in the preprocessing step to localize context and extract information.
@@ 6.2. Context Modeling Methods in Seven Code Intelligence Tasks
Rule-based (Jiang et al., 2007; Kamiya, 2015; Zhang et al., 2013; Zou et al., 2020; E. V. and Samuel, 2019; Koyuncu et al., 2019; Bowman and Huang, 2020; Huang et al., 2017; Cortés-Coy et al., 2014; Swe and Oo, 2019, 2019; Jiang et al., 2018; Wang and Lo, 2014; Wong et al., 2015; Motwani and Brun, 2023; Youm et al., 2017; McBurney and McMillan, 2016; Wen et al., 2018; Kim et al., 2019; Kim and Kim, 2019; Asaduzzaman et al., 2016, 2014; Saha et al., 2013; Nguyen et al., 2012; Wang et al., 2009; Zhang et al., 2023c; Shen et al., 2016; Yang et al., 2024b; Nafi et al., 2018; Zhang et al., 2024f) 29 29
@@
DL-based Sequence-based DNN (Lam et al., 2015; Nafi et al., 2019; Li et al., 2024f) 3
@@
LSTM (Lin et al., 2022; Lyu et al., 2021; Zhou et al., 2024; Chen et al., 2021a) 4
Bi-LSTM (Wang et al., 2021b; Yuan et al., 2020; Zhang et al., 2023d; Yu et al., 2023; Jiang et al., 2024) 5 45
LSTM-attention (Abhinav et al., 2021; Wang et al., 2021b; Zhou et al., 2019; Guo et al., 2023a) 4
GRU (Wang et al., 2021a; Bansal et al., 2021, 2023a; Liu et al., 2021; Yu et al., 2020; Zhang et al., 2021; Jiang et al., 2024) 7
Transformer (Li et al., 2023c; Washio and Miyao, 2022; Pei et al., 2023; Xu et al., 2022; Nie et al., 2021; Vu et al., 2023; Shahbazi et al., 2021; Ding et al., 2023b; Shrivastava et al., 2023a; Bansal et al., 2023c; Zügner et al., 2021; Liu et al., 2022; Guo et al., 2023b; Liu et al., 2023c; Wang et al., 2024f; Mandli et al., 2024; Lu et al., 2022; Jiang et al., 2024; Zhao et al., 2024; Wang et al., 2022; Li et al., 2024f) 21
@@
GCN (Nashid et al., 2023; Yuan et al., 2020; Mandli et al., 2024; Xu et al., 2021; Zhang et al., 2024a) 5 11
@@
LLM-based (Phan et al., 2024; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Wang et al., 2024g; Cheng et al., 2024; Tan et al., 2024; Li et al., 2024d; Wu et al., 2024b; Wang et al., 2024d; Liu et al., 2024b; Deng et al., 2024; Zhang et al., 2024b; Ma et al., 2024b; Li et al., 2024b; Bi et al., 2024b; Jiang et al., 2024; Eliseeva et al., 2023; Zhang et al., 2024h; Guan and Treude, 2024; Zhou et al., 2024; Ma et al., 2024a; Liu et al., 2024a; Li et al., 2024e; Chen et al., 2024b; Agrawal et al., 2024; Shrivastava et al., 2023b; Zhang et al., 2023a; Zan et al., 2022a; Chen et al., 2023; Zhou et al., 2023) 31 31
In code intelligence tasks, contextual data undergoes modeling after being preprocessed. However, traditional context-irrelevant approaches struggle to handle such contextual complexity effectively. This is due to the overwhelming scale of contextual data combined with sparse task-relevant information. Conventional methods struggle to capture the deep, hierarchical relationships and nuanced dependencies in complex repositories, failing to extract task-related information from context. These difficulties necessitate the development of new methods.
In this subsection, we focus on context-aware techniques presented in new technique papers, excluding benchmarks, case studies, and empirical studies. We identify 122 new technique papers on context-aware techniques and categorize their modeling approaches into four groups: rule, feature, deep learning, and large language models. Furthermore, if a paper implements a new context-aware method on models with more than 1 billion parameters, in that case, we classify these models as large language models rather than simply categorizing them as transformer-based.
@@ 6.2.1. Rule-Based
Rule-driven modeling methodologies, constructed based on domain-specific prior knowledge, exhibit advantages in specialized task scenarios. Rule-based models demonstrate two advantages. First, the verifiable nature of rule systems ensures output reliability. Second, rule-based architectures maintain predictable computational complexity.
As evidenced by 29 representative studies, this paradigm encodes expert knowledge into deterministic rules, well-suited for tasks such as program repair (Koyuncu et al., 2019; Jiang et al., 2018; Motwani and Brun, 2023; Wen et al., 2018; Kim et al., 2019; Kim and Kim, 2019) and defect detection (Swe and Oo, 2019; Wang and Lo, 2014; Youm et al., 2017; Zhao et al., 2011; Saha et al., 2013; Wang et al., 2009; Zhang et al., 2023c). For example, rule-based models facilitate coding intelligence tasks through predefined similarity assessment or matching protocols between existing context and target inputs. Wang et al. (Wang and Lo, 2014) exemplify this approach through three rule-governed similarity dimensions: version evolution patterns, historical reports correlations, and structural code correspondences. Their defect detection system applies verifiable scoring rules to generate prioritized suspicious file rankings through weighted aggregation. Jiang et al. (Jiang et al., 2018) leverage rule-driven AST context analysis to identify semantically code patterns relevant to historical patches. Their methodology systematically mines potential defective source code from existing code repositories, distilling recurrent repair templates through structural AST context rule matching.
@@ 6.2.3. Deep Learning-Based
Designed to capture long-term dependencies in sequential data, long short-term memory (LSTM) and gated recurrent unit (GRU) are two types of RNN introduced to solve code intelligence tasks (Abhinav et al., 2021; Wang et al., 2021b, a). LSTM uses gating mechanisms to regulate the dynamic retention and forgetting of information, preventing the vanishing or exploding gradient problems found in traditional RNNs. For example, Lin et al. (Zhang et al., 2023d) treat AST context as a sequence to input LSTM and predict patch correctness. Lyu et al. (Lyu et al., 2021) construct API dependency graphs as the input of LSTM for code generation. Compared with LSTM, GRU does not have a distinct memory cell but manages memory through its hidden state and improves computational efficiency by having a more straightforward gating mechanism. These advantages help the model maintain low memory usage and computational cost when utilizing context at scale. For example, some attempts (Haque et al., 2020; Bansal et al., 2021, 2023a) in code summarization employ GRU to hierarchically model context. Haque et al. (Haque et al., 2020) use one GRU per function to embed context function while utilizing a fully connected neural network to combine features from these RNNs and predict code summaries. Bansal et al. (Bansal et al., 2023a) employ similar methods by using one GRU to embed each called function and utilizing convGNN to combine embedding and summarize code. Bansal et al. (Bansal et al., 2021) construct a project-level embedding method for code summarization by constructing pipelines from word embedding, subroutine embedding, and file embedding to project embedding. Besides basic LSTM and GRU structure, researchers also employ bi-directional structure and attention mechanism to improve performance. For example, RepairNet (Abhinav et al., 2021) constructs an Attention-LSTM structure to receive compilation information cyclically for automatically compile-failed program repair. Yuan et al. (Yuan et al., 2020) trains a Bi-LSTM model to capture CFG features before computing feature similarity for clone detection. CoRec (Wang et al., 2021b) constructs a Bi-LSTM with an attention mechanism to model retrieved context and generate a commit message. Yu et al. (Yu et al., 2020) and Wang et al. (Wang et al., 2023) both use a Bi-GRU as part of the encoder and another GRU as a decoder for code comment generation. Yu et al. (Yu et al., 2020) utilize a GAT to capture graph representations between classes, and Wang et al. (Wang et al., 2023) employ a GAN to capture the relation between function graphs and comment graphs.
@@
Transformer (Vaswani et al., 2017) is one of the most popular models in recent years. Multi-head attention and self-attention mechanisms enable transformers to capture long-range dependency, which is suitable for processing long contextual information effectively. The most common way for a transformer is to add the contextual input type. For example, TransRepair (Li et al., 2023c) considers the inputs of diagnostic feedback from the compiler and source code context extracted by the lexical analyzer to repair the broken code. Washio and Miyao (Washio and Miyao, 2022) extract API document information and serve it as input for a transformer-based code generator. Many studies focus on precisely capturing relevant context before inputting the transformer, such as based on static analysis (Pei et al., 2023) or retrieval (Ding et al., 2023b; Lu et al., 2022). Other studies, however, focus on changing attention mechanisms. For example, LongCoder (Guo et al., 2023b) proposes window attention, global attention, and bridge attention to better capture source code context for code completion.
@@ 6.2.4. Large Language Model-Based
The emergence of large language models (LLM) changes the paradigm of context utilization in the code intelligence domain and shows a promising future in this area. Previous study often focuses on the construction of models (Zhang et al., 2023d; Haque et al., 2020; Bansal et al., 2021, 2023a; Koyuncu et al., 2019). However, researchers based on LLM often propose a framework to collect context and input it into LLMs (Phan et al., 2024; Wu et al., 2024b; Bi et al., 2024b; Li et al., 2024e). Usually, studies implement their methods on LLMs, such as StarCoder (Chen et al., 2023; Wu et al., 2024b; Liang et al., 2024; Wang et al., 2024d; Liu et al., 2024b; Deng et al., 2024; Liu et al., 2024a), Code-Llama (Bi et al., 2024b; Liang et al., 2024; Wang et al., 2024g; Cheng et al., 2024; Wang et al., 2024d; Deng et al., 2024; Li et al., 2024b; Liu et al., 2024a), CodeGen (Zhang et al., 2023a; Shahbazi and Fard, 2023; Li et al., 2024e), DeepSeek-Coder (Liang et al., 2024; Wang et al., 2024d; Deng et al., 2024; Liu et al., 2024a; Guan and Treude, 2024) and GPT-3.5-Turbo (Zhang et al., 2023a; Phan et al., 2024; Bi et al., 2024b; Chen et al., 2024b; Nashid et al., 2024; Liao et al., 2024; Cheng et al., 2024; Li et al., 2024d; Zhang et al., 2024b; Li et al., 2024b; Eliseeva et al., 2023; Guan and Treude, 2024; Ma et al., 2024a), and they use vanilla LLMs as a baseline to evaluate the effectiveness of the methods. These studies can be divided into three categories: retrieval-based (Zhou et al., 2023; Wu et al., 2024b; Qiu et al., 2024), static-analysis-based (Shrivastava et al., 2023b; Chen et al., 2024b; Nashid et al., 2024; Liu et al., 2024a), and iterative-based (Zhang et al., 2023a; Chen et al., 2023; Bi et al., 2024b; Li et al., 2024e).
@@
The basic pipeline of retrieval-based methods extracts relevant context from pre-built repositories. The context is merged with the task to form an augmented prompt that is subsequently input into a LLM to generate the answer. Among retrieval methods, for example, DocPrompting (Zhou et al., 2023) utilizes the BM25 algorithm to retrieve API documents. RepoFormer (Wu et al., 2024b) finetunes StarCoder (Li et al., 2023b) to determine whether the task requires retrieval. RepoHyper (Phan et al., 2024) constructs a repo-level semantic graph for global context representation and retrieval.
@@
Static analysis is another type of method specialized in collecting code-based context. By scrutinizing the code statically without compiling, methods based on static analysis extract context and build up prompts. Usually, static analysis supplies more fine-grained context compared with retrieval. For instance, RLPG (Shrivastava et al., 2023b) proposes ten context sources of prompt and constructs a classifier to determine which prompt source should be inputted to LLM based on the task. RLCE (Chen et al., 2024b) implements a static code analysis tool capable of automatically parsing the repository into code segments based on its structure and collects relevant context based on error information. LANCE (Nashid et al., 2024) proposes a lightweight and language-agnostic static analysis to extract context.
@@
Iterative-based methods include methods based on retrieval and static analysis though they are implemented through an iterative pipeline. To utilize iterative-based frameworks, RepoCoder (Zhang et al., 2023a) combines the generation based on retrieval to further iterative retrieve and generate. RepoMinCoder (Li et al., 2024e), based on RepoCoder, adds an information-loss-based method using minimum description length to select and rerank retrieved context. CompVPD (Chen et al., 2023) introduces iterative human validation to generate a precise context. CocoGen (Bi et al., 2024b) iterative retrieves context based on compiler feedback.
@@
☞ RQ3 Summary: (1) We summarize the data preprocessing steps discussed in the reviewed papers. The most commonly used method is to split the identifiers and perform text normalization on the source code. (2) We analyze the structure of context-aware models, finding that sequence-based deep learning models are the most widely used. Additionally, researchers prefer employing tree-based and graph-based models to process specific types of context, such as abstract syntax trees (ASTs) and control flow graphs (CFGs). (3) There is a trend (31 studies) exploring the usage of LLMs to process context, presenting significant opportunities for further research in this area.

### v2

@@ 6.1.2. Removal
For natural language data, stopwords are often removed due to their high frequency and little meaningful contribution (e.g., ”a” and ”the”) (Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017). For code-formatted data, heuristic rules can be designed to clean the data effectively. Irrelevant elements, such as logical operators (Matmti and Fard, 2023), trailing white spaces (Matmti and Fard, 2023), punctuation marks (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021b; Bansal et al., 2023a; Koyuncu et al., 2019; Matmti and Fard, 2023; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017), comments (Zügner et al., 2021), empty lines (Zügner et al., 2021), and non-word characters (Bansal et al., 2023a), can also be removed to simplify the data and reduce noise.
@@ 6.1.3. Unifying
To organize natural language formatted data, common text unifying methods include using lowercase (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021b) and applying stemming (Koyuncu et al., 2019; Bangcharoensap et al., 2012; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017). These works simplify the data by standardizing variations in word form. Lowercasing ensures that all tokens are converted to lowercase while stemming groups of words with the same core meaning by retaining the root form of words. Both methods help models focus on the core meaning of the data rather than extra information.
@@
In addition to organizing data, handling sensitive information is crucial to ensure data purity and model security. Researchers replace numbers and strings in source code with unique tokens such as ¡num¿ (Nashid et al., 2023; Xu et al., 2022; Wang et al., 2021b; Zügner et al., 2021; Lin et al., 2021) and ¡str¿ (Nashid et al., 2023; Xu et al., 2022; Zhou et al., 2019; Wang et al., 2021b; Bansal et al., 2023a; Zügner et al., 2021; Lin et al., 2021). Numbers and strings can vary in length and format, requiring additional handling methods. Masking specific values encourages models to focus on the data type. Additionally, replacing them with unique tokens helps protect private data, since numbers and strings may sometimes contain sensitive information.
@@ 6.1.4. Other
One of the preprocessing methods involves using ASTs to extract information from source code. In this case, ASTs are not considered part of the context but are used to extract semantic code knowledge from repositories (Bi et al., 2024; Liao et al., 2024). This narrows the scope of source code context while facilitating retrieval by constructing semantic graphs (Phan et al., 2024).
Besides, researchers also use static analysis to convert source code into XML format (Zhou et al., 2024) or apply program slicing (Zhang et al., 2024d), enabling further analysis and processing. In addition, manual expertise (Zhao et al., 2024) and rule-based methods (Wang et al., 2023) are also applied in the preprocessing step to localize context and extract information.
@@ 6.2. Context Modeling Methods in Seven Code Intelligence Tasks
Rule-based (Jiang et al., 2007; Kamiya, 2015; Zhang et al., 2013; Zou et al., 2020; E. V. and Samuel, 2019; Koyuncu et al., 2019; Bowman and Huang, 2020; Huang et al., 2017; Cortés-Coy et al., 2014; Swe and Oo, 2019; Jiang et al., 2018; Wang and Lo, 2014; Wong et al., 2015; Motwani and Brun, 2023; Youm et al., 2017; McBurney and McMillan, 2016; Wen et al., 2018; Kim et al., 2019; Kim and Kim, 2019; Asaduzzaman et al., 2016, 2014; Saha et al., 2013; Nguyen et al., 2012; Wang et al., 2009; Zhang et al., 2023d; Shen et al., 2016; Yang et al., 2024; Nafi et al., 2018; Zhang et al., 2024c) 29 29
@@
DL-based Sequence-based DNN (Lam et al., 2015; Nafi et al., 2019; Li et al., 2024b) 3
@@
LSTM (Lin et al., 2022; Lyu et al., 2021; Zhou et al., 2024; Chen et al., 2021b) 4
Bi-LSTM (Wang et al., 2021a; Yuan et al., 2020; Zhang et al., 2023c; Yu et al., 2023; Jiang et al., 2024) 5 45
LSTM-attention (Abhinav et al., 2021; Wang et al., 2021a; Zhou et al., 2019; Guo et al., 2023b) 4
GRU (Wang et al., 2021b; Bansal et al., 2021, 2023a; Liu et al., 2021; Yu et al., 2020; Zhang et al., 2021; Jiang et al., 2024) 7
Transformer (Li et al., 2023b; Washio and Miyao, 2022; Pei et al., 2023; Xu et al., 2022; Nie et al., 2021; Vu et al., 2023; Shahbazi et al., 2021; Ding et al., 2023b; Shrivastava et al., 2023a; Bansal et al., 2023c; Zügner et al., 2021; Liu et al., 2022; Guo et al., 2023a; Liu et al., 2023b; Wang et al., 2024b; Mandli et al., 2024; Lu et al., 2022; Jiang et al., 2024; Zhao et al., 2024; Wang et al., 2022; Li et al., 2024b) 21
@@
GCN (Nashid et al., 2023; Yuan et al., 2020; Mandli et al., 2024; Xu et al., 2021; Zhang et al., 2024d) 5 11
@@
LLM-based (Phan et al., 2024; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Wang et al., 2024a; Cheng et al., 2024; Tan et al., 2024; Li et al., 2024e; Wu et al., 2024a; Wang et al., 2024g; Liu et al., 2024b; Deng et al., 2024; Zhang et al., 2024a; Ma et al., 2024a; Li et al., 2024a; Bi et al., 2024; Jiang et al., 2024; Eliseeva et al., 2023; Zhang et al., 2024b; Guan and Treude, 2024; Zhou et al., 2024; Ma et al., 2024b; Liu et al., 2024a; Li et al., 2024f; Chen et al., 2024; Agrawal et al., 2024; Zan et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Zan et al., 2022; Chen et al., 2023; Zhou et al., 2023) 32 32
In code intelligence tasks, contextual data is modeled after preprocessing. However, traditional context-irrelevant approaches struggle to handle such contextual complexity effectively. This is due to the overwhelming scale of contextual data combined with sparse task-relevant information. Conventional methods struggle to capture the deep, hierarchical relationships and nuanced dependencies in complex repositories, failing to extract task-related information from context. These difficulties necessitate developing new methods.
In this subsection, we focus on context-aware techniques presented in new technique papers, excluding benchmarks, case studies, and empirical studies. We identify 122 new technique papers on context-aware techniques and categorize their modeling approaches into four groups: rule, feature, deep learning, and large language models. Furthermore, if a paper implements a new context-aware method on models with more than 1 billion parameters, we classify these models as large language models rather than simply as transformer-based.
@@ 6.2.1. Rule-Based
Rule-driven modeling methodologies, constructed from domain-specific prior knowledge, have advantages in specialized task scenarios. Rule-based models demonstrate two advantages. First, the verifiable nature of rule systems ensures output reliability. Second, rule-based architectures maintain predictable computational complexity.
As evidenced by 29 representative studies, this paradigm encodes expert knowledge into deterministic rules, well-suited for tasks such as program repair (Koyuncu et al., 2019; Jiang et al., 2018; Motwani and Brun, 2023; Wen et al., 2018; Kim et al., 2019; Kim and Kim, 2019) and defect detection (Swe and Oo, 2019; Wang and Lo, 2014; Youm et al., 2017; Zhao et al., 2011; Saha et al., 2013; Wang et al., 2009; Zhang et al., 2023d). For example, rule-based models facilitate coding intelligence tasks through predefined similarity assessment or matching protocols between existing context and target inputs. Wang et al. (Wang and Lo, 2014) exemplify this approach through three rule-governed similarity dimensions: version evolution patterns, historical reports correlations, and structural code correspondences. Their defect detection system applies verifiable scoring rules to generate prioritized suspicious file rankings through weighted aggregation. Jiang et al. (Jiang et al., 2018) leverage rule-driven AST context analysis to identify semantically code patterns relevant to historical patches. Their methodology systematically mines potential defective source code from existing code repositories, distilling recurrent repair templates through structural AST context rule matching.
@@ 6.2.3. Deep Learning-Based
Designed to capture long-term dependencies in sequential data, long short-term memory (LSTM) and gated recurrent unit (GRU) are two types of RNN introduced to solve code intelligence tasks (Abhinav et al., 2021; Wang et al., 2021a, b). LSTM uses gating mechanisms to regulate the dynamic retention and forgetting of information, preventing the vanishing or exploding gradient problems found in traditional RNNs. For example, Lin et al. (Zhang et al., 2023c) treat AST context as a sequence to input LSTM and predict patch correctness. Lyu et al. (Lyu et al., 2021) construct API dependency graphs as the input of LSTM for code generation. Compared with LSTM, GRU does not have a distinct memory cell but manages memory through its hidden state and improves computational efficiency by having a more straightforward gating mechanism. These advantages help the model maintain low memory usage and computational cost when utilizing context at scale. For example, some attempts (Haque et al., 2020; Bansal et al., 2021, 2023a) in code summarization employ GRU to hierarchically model context. Haque et al. (Haque et al., 2020) use one GRU per function to embed context function while utilizing a fully connected neural network to combine features from these RNNs and predict code summaries. Bansal et al. (Bansal et al., 2023a) employ similar methods by using one GRU to embed each called function and utilizing convGNN to combine the embedding and summarize the code. Bansal et al. (Bansal et al., 2021) construct a project-level embedding method for code summarization by constructing pipelines from word embedding, subroutine embedding, and file embedding to project embedding. Besides basic LSTM and GRU structures, researchers also employ bi-directional structures and attention mechanisms to improve performance. For example, RepairNet (Abhinav et al., 2021) constructs an Attention-LSTM structure to receive compilation information cyclically for automatically compiling failed program repair. Yuan et al. (Yuan et al., 2020) trains a Bi-LSTM model to capture CFG features before computing feature similarity for clone detection. CoRec (Wang et al., 2021a) constructs a Bi-LSTM with an attention mechanism to model retrieved context and generate a commit message. Yu et al. (Yu et al., 2020) and Wang et al. (Wang et al., 2023) both use a Bi-GRU as part of the encoder and another GRU as a decoder for code comment generation. Yu et al. (Yu et al., 2020) utilize a GAT to capture graph representations between classes, and Wang et al. (Wang et al., 2023) employ a GAN to capture the relation between function graphs and comment graphs.
@@
Transformer (Vaswani et al., 2017) is one of the most popular models in recent years. Multi-head attention and self-attention mechanisms enable transformers to capture long-range dependencies, making them well-suited to processing long contextual information effectively. The most common way for a transformer is to add the contextual input type. For example, TransRepair (Li et al., 2023b) considers the inputs of diagnostic feedback from the compiler and source code context extracted by the lexical analyzer to repair the broken code. Washio and Miyao (Washio and Miyao, 2022) extract API document information and serve it as input for a transformer-based code generator. Many studies focus on precisely capturing relevant context before inputting the transformer, such as based on static analysis (Pei et al., 2023) or retrieval (Ding et al., 2023b; Lu et al., 2022). Other studies, however, focus on changing attention mechanisms. For example, LongCoder (Guo et al., 2023a) proposes window attention, global attention, and bridge attention to better capture source code context for code completion.
@@ 6.2.4. Large Language Model-Based
The emergence of LLMs changes the paradigm of context utilization in the code intelligence domain and shows a promising future in this area. Crucially, LLM-based methods represent a fundamental shift in context handling compared to earlier deep learning architectures. The primary contrast lies in the mechanism of context integration.
@@
Prior deep learning models integrate context through architectural modifications explicitly designed to model contexts (Zhang et al., 2023c; Haque et al., 2020; Bansal et al., 2021, 2023a; Koyuncu et al., 2019). This approach is highly effective but demands specialized engineering for each context type. In contrast, LLM-based methods leverage context augmentation via prompt Engineering and retrieval (Phan et al., 2024; Wu et al., 2024a; Bi et al., 2024; Li et al., 2024f). They convert diverse context types into a sequential augmented prompt. This shift means that the focus moves from modeling the context structure (DL-based) to retrieving and formulating the most relevant context (LLM-based). Usually, studies implement their methods on LLMs, such as StarCoder (Chen et al., 2023; Wu et al., 2024a; Liang et al., 2024; Wang et al., 2024g; Liu et al., 2024b; Deng et al., 2024; Liu et al., 2024a), Code-Llama (Bi et al., 2024; Liang et al., 2024; Wang et al., 2024a; Cheng et al., 2024; Wang et al., 2024g; Deng et al., 2024; Li et al., 2024a; Liu et al., 2024a), CodeGen (Zhang et al., 2023a; Shahbazi and Fard, 2023; Li et al., 2024f), DeepSeek-Coder (Liang et al., 2024; Wang et al., 2024g; Deng et al., 2024; Liu et al., 2024a; Guan and Treude, 2024) and GPT-3.5-Turbo (Zhang et al., 2023a; Phan et al., 2024; Bi et al., 2024; Chen et al., 2024; Nashid et al., 2024; Liao et al., 2024; Cheng et al., 2024; Li et al., 2024e; Zhang et al., 2024a; Li et al., 2024a; Eliseeva et al., 2023; Guan and Treude, 2024; Ma et al., 2024b), and they use vanilla LLMs as a baseline to evaluate the effectiveness of the methods. These studies can be divided into three categories: retrieval-based (Zhou et al., 2023; Wu et al., 2024a; Qiu et al., 2024), static-analysis-based (Shrivastava et al., 2023b; Chen et al., 2024; Nashid et al., 2024; Liu et al., 2024a), and iterative-based (Zhang et al., 2023a; Chen et al., 2023; Bi et al., 2024; Li et al., 2024f).
@@
The basic pipeline of retrieval-based methods extracts relevant context from pre-built repositories. The context is merged with the task to form an augmented prompt that is subsequently input into a LLM to generate the answer. Among retrieval methods, for example, DocPrompting (Zhou et al., 2023) utilizes the BM25 algorithm to retrieve API documents. RepoFormer (Wu et al., 2024a) finetunes StarCoder (Li et al., 2023a) to determine whether the task requires retrieval. RepoHyper (Phan et al., 2024) constructs a repo-level semantic graph for global context representation and retrieval.
@@
Static analysis is another type of method specialized in collecting code-based context. By scrutinizing the code statically without compiling, methods based on static analysis extract context and build up prompts. Usually, static analysis supplies more fine-grained context compared with retrieval. For instance, RLPG (Shrivastava et al., 2023b) proposes ten context sources of prompt and constructs a classifier to determine which prompt source should be inputted to LLM based on the task. RLCE (Chen et al., 2024) implements a static code analysis tool capable of automatically parsing the repository into code segments based on its structure and collects relevant context based on error information. LANCE (Nashid et al., 2024) proposes a lightweight and language-agnostic static analysis to extract context.
@@
Iterative-based methods include methods based on retrieval and static analysis, though they are implemented through an iterative pipeline. To utilize iterative-based frameworks, RepoCoder (Zhang et al., 2023a) combines the generation based on retrieval to further iteratively retrieve and generate. RepoMinCoder (Li et al., 2024f), based on RepoCoder, adds an information-loss-based method using minimum description length to select and rerank retrieved context. CompVPD (Chen et al., 2023) introduces iterative human validation to generate a precise context. CocoGen (Bi et al., 2024) iteratively retrieves context based on compiler feedback.
@@
☞RQ3 Summary: (1) We summarize the data preprocessing steps discussed in the reviewed papers. The most commonly used method is to split the identifiers and perform text normalization on the source code. (2) We analyze the structure of context-aware models, finding that sequence-based deep learning models are the most widely used. Additionally, researchers prefer employing tree-based and graph-based models to process specific types of context, such as abstract syntax trees (ASTs) and control flow graphs (CFGs). (3) There is a trend (32 studies) exploring the usage of LLMs to process context, presenting significant opportunities for further research in this area.